OpenStudy (anonymous):

"The length of a rectangular playground is 25 meters more than three times the width. It took 650 meters of fencing to enclose it. Find the dimensions of the playground." Anyone that could help would be amazing!!

OpenStudy (anonymous):

length L=25+3W perimeter P=650=2*(L+W) solve the problem by substituting value of L in second equation... try it :)

OpenStudy (anonymous):

P=2(25+3w+w)=650 25+4w=325 4w=300 w=75 so you know your width, now: L=25+3*75=250 just to verify your answer - check - P=650 :)